Accountancy

The goodwill of a firm is estimated at three years purchase of the average profits of the last 5 years are as follows 

Year   profit 

1999  10000

2000    15000

2001    5000

2002     5000loss

2003    8000
Answer #1

Average profit=10000+15000+5000+8000-5000/5

=33000/5=6600

Goodwill=6600×3=19800

answered by: Advika